OpenStudy (anonymous):

graph the equation -3x-4y=12

OpenStudy (ibbutibbu.):

i would put it into the form y = mx + b which is y = -3/4x - 3

OpenStudy (ibbutibbu.):

then graph the point (0,-3)

OpenStudy (ibbutibbu.):

then go down 3 units, and to the right 4 and you have oyour line!

OpenStudy (ibbutibbu.):

does that clear it up for ypu, @penfold

OpenStudy (anonymous):

-3x-4y=12 -4y=12+3x y=12+3x/-4